OpenStudy (anonymous):

How do I calculate the following: (sqrt(3) * sqrt(3) * sqrt(3)) / (sqrt(3) + sqrt(3) + sqrt(3)) Answer is : 1

OpenStudy (anonymous):

\[(\sqrt{3} * \sqrt{3} * \sqrt{3}) / (\sqrt{3} + \sqrt{3} + \sqrt{3})\]

OpenStudy (madhatter):

here this might help https://www.youtube.com/watch?v=K067tun3UEo

OpenStudy (anonymous):

hmm, not really :/

OpenStudy (madhatter):

oh sorry thought it would

OpenStudy (anonymous):

\[\sqrt{3} * \sqrt{3} * \sqrt{3} = \sqrt{3*3*3}\]

OpenStudy (anonymous):

\[\sqrt{4} + \sqrt{4} = 2 + 2 = 4\] Simple. But seeing as how that's the level we're at, I assuming there's some convenient shortcut for \[\sqrt{3} + \sqrt{3}\]

OpenStudy (anonymous):

both numerator and denominator are \(3\sqrt3\)

OpenStudy (anonymous):

sorry my drawing is bad but you would do the sqrt of 3 which is 1.7 then, 1.7x1.7x1.7 = 4.9 then, 1.7+1.7+1.7 or 1.7x3 =5.1 then, 5.1/4.9

OpenStudy (anonymous):

\[\sqrt{3}\sqrt{3}\sqrt{3}=3\sqrt{3}\] and \[\sqrt3+\sqrt3+\sqrt3=3\sqrt3\]

OpenStudy (anonymous):

you get \[\frac{3\sqrt3}{3\sqrt3}=1\]

OpenStudy (anonymous):

i would do it my way

OpenStudy (anonymous):

Interesting... how come satellites technique doesn't work for \[\sqrt{5} * \sqrt{5}\] and \[\sqrt{5} + \sqrt{5}\] ?

OpenStudy (anonymous):

works fine for the problem I presented